# Stability and bifurcation analysis in a mechanochemical model of pattern formation

### v1

@@ 6.3. Global continuation and turning points
Appendix A Well-posedness of the model
@@
(A.1) $u(t,\cdot)=e^{D\Delta t}u_{0}+\int_{0}^{t}e^{D\Delta(t-s)}\left(\kappa\frac{e^{u(s,\cdot)}}{\int_{0}^{1}e^{u(s,y)}dy}-u(s,\cdot)\right)ds\;.$
@@
(A.2) $F(v)=e^{D\Delta t}u_{0}+\int_{0}^{t}e^{D\Delta(t-s)}\left(\kappa\frac{e^{v(s,\cdot)}}{\int_{0}^{1}e^{v(s,y)}dy}-v(s,\cdot)\right)ds.$
@@
(A.3) $V=\left\{v\in X:\|v\|_{X}\leqslant 3\delta\right\}\quad\text{where}\quad\delta=\|u_{0}\|_{W^{1,2}}.$
@@
(A.4) $\|e^{D\Delta t}v\|_{L^{2}}\leqslant C\|v\|_{L^{2}}\quad\text{and}\quad\|\nabla e^{D\Delta t}v\|_{L^{2}}\leqslant Ct^{-\frac{1}{2}}\|v\|_{L^{2}}\quad\text{for }v\in L^{2}(\mathbb{T}).$
@@
(A.5) $\|F(v)\|_{W^{1,2}}$ $\leqslant\|e^{D\Delta t}u_{0}\|_{W^{1,2}}+C\int_{0}^{t}\left\|e^{D\Delta(t-s)}\left(\kappa\frac{e^{v(s,\cdot)}}{\int_{0}^{1}e^{v(s,y)}dy}-v(s,\cdot)\right)\right\|_{W^{1,2}}ds$
@@
(A.6) $\|F(v)\|_{W^{1,2}}$ $\leqslant\|e^{D\Delta t}u_{0}\|_{W^{1,2}}+C\int_{0}^{t}\left(1+(t-s)^{-\frac{1}{2}}\right)\left(\kappa\frac{e^{\|v(s,\cdot)\|_{L^{\infty}}}}{e^{-\|v(s,\cdot)\|_{L^{\infty}}}}+\|v(s,\cdot)\|_{2}\right)ds$
@@
(A.7) $\|F(v)$ $-F(w)\|_{W^{1,2}}\leqslant\int_{0}^{t}\left(1+(t-s)^{-\frac{1}{2}}\right)\left\|\kappa\frac{e^{v}}{\int_{0}^{1}e^{v}dy}-\kappa\frac{e^{w}}{\int_{0}^{1}e^{w}dy}-v+w\right\|_{L^{2}}ds.$
@@
(A.8) $\|F(v)-F(w)\|_{W^{1,2}}$ $\leqslant C(t)\|v-w\|_{X}+C(t)\kappa\sup_{t\in[0,\tau]}\left\|\kappa\frac{e^{v}-e^{w}}{\int_{0}^{1}e^{v}dy}+\kappa e^{w}\frac{\int_{0}^{1}e^{w}dy-\int_{0}^{1}e^{v}dy}{\int_{0}^{1}e^{v}dy\int_{0}^{1}e^{w}dy}\right\|_{L^{2}}.$
@@
(A.9) $e^{w}\left|\int_{0}^{1}(e^{w}-e^{v})dy\right|\big/\left(\int_{0}^{1}e^{v}dy\int_{0}^{1}e^{w}dy\right)\leqslant e^{3C\delta}\int_{0}^{1}|e^{v}-e^{w}|dy$
@@
(A.11) $\|F(v)-F(w)\|_{W^{1,2}}$ $\leqslant C(t)\|v-w\|_{X}+C(t)\kappa e^{C\delta}\sup_{t\in[0,\tau]}\left\|{e^{v}-e^{w}}\right\|_{L^{2}}+C(t)\kappa e^{3C\delta}\sup_{t\in[0,\tau]}\left\|{e^{v}-e^{w}}\right\|_{L^{\infty}}.$
@@
(A.12) $\|F(v)-F(w)\|_{W^{1,2}}\leqslant C(t)\|v-w\|_{X}+C(t)\sup_{t\in[0,\tau]}\left\|{v-w}\right\|_{W^{1,2}}\leqslant C(\tau)\|v-w\|_{X}.$
@@
(A.13) $\frac{d}{dt}\mathcal{J}\big(u(\cdot,t)\big)=D\mathcal{J}\big(u(\cdot,t)\big)u_{t}(\cdot,t)=-\|u_{t}(\cdot,t)\|_{L^{2}}^{2}\leqslant 0\;,$
@@
(A.14) $\frac{D}{2}\int_{0}^{1}\big|u_{x}(x,t)\big|^{2}dx+\frac{1}{2}\int_{0}^{1}\big|u(x,t)\big|^{2}dx$ $\leqslant\mathcal{J}\big(u(\cdot,t)\big)+\kappa\log\left(\int_{0}^{1}e^{u(x,t)}dx\right)$
@@
(A.15) $\int_{0}^{t}\|u_{t}(\cdot,t)\|_{L^{2}}^{2}$ $=J(u_{0})-J\big(u(\cdot,t)\big)\leqslant J(u_{0})+C\|u(\cdot,t)\|_{W{1,2}}+\kappa C\|u(\cdot,t)\|_{L^{\infty}}\leqslant C\;,$
@@
(A.16) $\int_{0}^{t}\|\Delta u(\cdot,t)\|_{L^{2}}^{2}$ $=\int_{0}^{t}\left\|u_{t}(\cdot,t)+u(\cdot,t)+\kappa\frac{e^{u(\cdot,t)}}{\int_{0}^{1}e^{u(y,t)}dy}\right\|_{L^{2}}^{2}$
@@
(A.17) $u\in C\big([0,\infty),W^{1,2}(\mathbb{T})\big),\qquad u\in L^{2}\big([0,\infty),L^{2}(\mathbb{T})\big),\qquad u\in L^{2}\big((0,T),W^{2,2}(\mathbb{T})\big),$
@@
Appendix B Computations of the derivatives
@@
(B.1) $G_{\chi}(\chi,\alpha)(h)$ $=Dh_{xx}-h+(\kappa_{n}+\alpha)\left[\frac{e^{\chi}}{\int e^{\chi}dy}h-\frac{e^{\chi}}{(\int e^{\chi}dy)^{2}}\int e^{\chi}hdy\right],$

### v2

@@ 6.3. Global continuation and turning points
Appendix A Model derivation
Figure A.1. Schematic representation of the elastic deformation framework for the Hydra spheroid model. Left: Reference configuration $\mathcal{S}_{0}$ parametrized by $\mathbf{g}:[0,L]\to\mathcal{S}_{0}$. Right: Deformed configuration $\mathcal{S}$ showing the composition of mappings $\mathbf{f}=\mathbf{h}\circ\mathbf{g}$ and the arc-length reparametrization $\mathbf{\tilde{f}}$ with $\mathbf{f}=\mathbf{\tilde{f}}\circ s$.
A.1. Hyper-elastic tissue model
We demonstrate the derivation of the mechanochemical patterning model for Hydra spheroids introduced in [41] in two spatial dimensions. It retains the essential coupling between mechanics and chemistry while simplifying the geometric structure. In this setting, the Hydra spheroid is represented as an elastic closed curve, providing a minimal framework to analyze the core mechanochemical feedback mechanisms, see Figure A.1.
(A.1) $\mathbf{g}(\xi)=\frac{L}{2\pi}\begin{pmatrix}\cos\frac{2\pi}{L}\xi\\ \sin\frac{2\pi}{L}\xi\end{pmatrix},\quad\xi\in[0,L].$
This reference curve is subjected to a spatial deformation $\mathbf{h}:\mathcal{S}_{0}\to\mathcal{S}$, resulting in a new closed curve $\mathcal{S}$ with circumference $\ell>0$. We prescribe a fixed enclosed area $A_{0}>\operatorname{Area}(\mathcal{S}_{0})$ and restrict attention to deformations that satisfy $\operatorname{Area}(\mathcal{S})=A_{0}$. The deformed curve is parametrized by the composition
(A.2) $\mathbf{f}:[0,L]\to\mathcal{S},\quad\mathbf{f}=\mathbf{h}\circ\mathbf{g}.$
The deformation gradient in surface coordinates is given by the matrix
(A.3) $\mathbf{F}=\nabla_{\mathbf{X}}\mathbf{h}(\mathbf{X})=\mathbf{f}^{\prime}(\xi)\otimes\nabla\big(\mathbf{g}^{-1}(\mathbf{X})\big),\quad\mathbf{X}=\mathbf{g}(\xi)\;.$
Taking a derivative of the identity $\mathbf{g}^{-1}(\mathbf{g}(\xi))=\xi$ and using $|\mathbf{g}^{\prime}(\xi)|\equiv 1$ implies that $\nabla(\mathbf{g}^{-1}(\mathbf{X}))=\mathbf{g}^{\prime}(\xi)^{T}$, where $\mathbf{g}^{\prime}(\xi)$ (the covariant basis vector on $\mathcal{S}_{0}$) is a column vector and the contravariant basis vector $\mathbf{g}^{\prime}(\xi)^{T}$ is a row vector. We therefore can write
(A.4) $\mathbf{F}=\mathbf{f}^{\prime}(\xi)\otimes\mathbf{g}^{\prime}(\xi)^{T}.$
The Right Cauchy-Green deformation tensor reads
(A.5) $\mathbf{C}=\mathbf{F}^{T}\mathbf{F}=\left(\mathbf{g}^{\prime}(\xi)\otimes\mathbf{f}^{\prime}(\xi)^{T}\right)\left(\mathbf{f}^{\prime}(\xi)\otimes\mathbf{g}^{\prime}(\xi)^{T}\right)=|\mathbf{f}^{\prime}(\xi)|^{2}\,\mathbf{g}^{\prime}(\xi)\otimes\mathbf{g}^{\prime}(\xi)^{T}\;,$
and the Green-Lagrangian strain is
(A.6) $\boldsymbol{\varepsilon}=\frac{1}{2}(\mathbf{C}-\mathbf{I})=\frac{1}{2}\left(|\mathbf{f}^{\prime}(\xi)|^{2}-1\right)\mathbf{g}^{\prime}(\xi)\otimes\mathbf{g}^{\prime}(\xi)^{T},$
where the identity in the tangent space of $S_{0}$ is given by $\mathbf{I}=\mathbf{g}^{\prime}(\xi)\otimes\mathbf{g}^{\prime}(\xi)^{T}$.
(A.7) $W(\boldsymbol{\varepsilon})=\frac{\lambda}{2}{\rm tr}^{2}(\boldsymbol{\varepsilon})+\frac{\mu}{2}{\rm tr}(\boldsymbol{\varepsilon}^{2}),$
where the Lamé parameters satisfy $\lambda=0$ and $\mu=E(u)$, with $u:=u(\xi)$ the morphogen concentration and $E(u)$ the morphogen-dependent elastic modulus encoding the mechanochemical coupling central to our model. The specific choice of parameters follows from the modelling considerations in [41]. The total elastic energy of a deformation is given by
(A.8) $\mathcal{E}_{\text{el}}=\int_{S_{0}}W(\varepsilon)\,dS_{0}=\frac{1}{8}\int_{0}^{L}{E(u)}\left(|\mathbf{f}^{\prime}(\xi)|^{2}-1\right)^{2}|\mathbf{g}^{\prime}(\xi)|d\xi,$
where we used equation (A.6).
(A.9) $A_{0}=A(f)=\frac{1}{2}\int_{S}\mathbf{x}\cdot\mathbf{n}\,dS=\frac{1}{2}\int_{0}^{L}\mathbf{f}(\xi)\cdot\mathbf{n}(\mathbf{f}(\xi))\,|\mathbf{f}^{\prime}|\ {d}\xi=-\frac{1}{2}\int_{0}^{L}\mathbf{f}(\xi)\cdot\mathbf{f}^{\prime}(\xi)^{\perp}\ {d}\xi\;,$
where $\mathbf{n}(\mathbf{x})=-\mathbf{f}^{\prime}(\mathbf{x})^{\perp}/|\mathbf{f}^{\prime}(\mathbf{x})|$ is the outer unit normal vector on $\mathcal{S}$ at $\mathbf{x}\in\mathcal{S}$ and represents the rotation of a 2-dimensional vector by $\pi/2$: $(a,b)^{\perp}=(-b,a)$.
To find minimizers of $\mathcal{E}_{\text{el}}$ we introduce Lagrange multipliers. The Gateaux derivative with respect to variations $\delta\mathbf{f}(\xi)$ with periodic boundary conditions yields
(A.10) $0$ $=D\mathcal{E}_{\text{el}}(\mathbf{f})(\delta\mathbf{f})+pDA(f)(\delta f)$
$=\frac{1}{2}\int_{0}^{L}E(u)\left(|\mathbf{f}^{\prime}(\xi)|^{2}-1\right)\mathbf{f}^{\prime}(\xi)\cdot\delta\mathbf{f}^{\prime}(\xi)d\xi-p\int_{0}^{L}\delta\mathbf{f}(\xi)\cdot\mathbf{f}^{\prime}(\xi)^{\perp}\,d\xi\;,$
where $p\in\mathbb{R}$, the inner pressure, acts as a Lagrange multiplier for the area constraint. Integrating by parts, we obtain that equation (A.10) is satisfied if and only if the minimizer $f$ satisfies
(A.11) $\left(\sigma(\xi)\frac{\mathbf{f}^{\prime}(\xi)}{|\mathbf{f}^{\prime}(\xi)|}\right)^{\prime}+p\,\mathbf{f}^{\prime}(\xi)^{\perp}=0\;,$
where the stress is given by
(A.12) $\sigma(\xi)=E(u)\frac{1}{2}\left(|\mathbf{f}^{\prime}(\xi)|^{2}-1\right)|\mathbf{f}^{\prime}(\xi)|.$
Evaluating the derivative we find that
(A.13) $\sigma^{\prime}(\xi)\frac{\mathbf{f}^{\prime}(\xi)}{|\mathbf{f}^{\prime}(\xi)|}+\sigma(\xi)\left(\frac{\mathbf{f}^{\prime}(\xi)}{|\mathbf{f}^{\prime}(\xi)|}\right)^{\prime}+p\,\mathbf{f}^{\prime}(\xi)^{\perp}=0\;.$
Taking the scalar product of this with $\mathbf{f}^{\prime}(\xi)$ and using that $\left(\frac{\mathbf{f}^{\prime}(\xi)}{|\mathbf{f}^{\prime}(\xi)|}\right)^{\prime}\cdot\mathbf{f}^{\prime}(\xi)=0$ as well as the impenetrability assumption $|\mathbf{f}^{\prime}(\xi)|>0$ we obtain that $\sigma^{\prime}\equiv 0$, i.e. the stress is constant. Taking the scalar product of (A.13) with $\mathbf{f}^{\prime}(\xi)^{\perp}$ we obtain that $\sigma\left(\frac{\mathbf{f}^{\prime}(\xi)}{|\mathbf{f}^{\prime}(\xi)|}\right)^{\prime}\cdot\mathbf{f}^{\prime}(\xi)^{\perp}+p\,|\mathbf{f}^{\prime}(\xi)|^{2}=0$, i.e. the curvature $\kappa=\frac{\mathbf{f}^{\prime\prime}(\xi)}{|\mathbf{f}^{\prime}(\xi)|^{3}}\cdot\mathbf{f}^{\prime}(\xi)^{\perp}=-p/\sigma$ (Laplace’s Law) is constant along $\mathcal{S}$. Hence, the minimizer $\mathcal{S}$ must be circular with circumference $\ell>L$.
(A.14) $\mathbf{f}(\xi)=\mathbf{\tilde{f}}(s(\xi)),$
where $\mathbf{\tilde{f}}:[0,\ell]\to\mathcal{S}$ satisfies $|\mathbf{\tilde{f}}^{\prime}|=1$, and $s:[0,L]\to[0,\ell]$ with $s^{\prime}>0$ encodes the tangential deformation. Since $s^{\prime}(\xi)=|f^{\prime}(\xi)|$ represents the local stretch ratio (ratio of deformed to reference arc length $|g^{\prime}({\xi})|\equiv 1$), the quantity
(A.15) $\tilde{\varepsilon}=\frac{1}{2}((s^{\prime})^{2}-1)$
is the finite Green-Lagrangian strain in one dimension.
Conservation of total length imposes
(A.16) $\int_{0}^{L}(s^{\prime}(\xi)-1)\,d\xi=\ell-L.$
The morphogen concentration $u(t,\xi)$ evolves according to
(A.17) $\partial_{t}u=\kappa\tilde{\varepsilon}-\alpha u+D\partial_{\xi\xi}u,$
with production proportional to local strain, linear degradation rate $\alpha>0$, and diffusion coefficient $D>0$. Taking together, the full mechanochemical system reads
(A.18) $\left\{\begin{aligned} &\begin{aligned} \partial_{t}u&=\kappa\tilde{\varepsilon}-\alpha u+D\partial_{\xi\xi}u,\\ 0&=\partial_{\xi}\big(E(u)(s^{\prime 2}-1)s^{\prime}\big),\end{aligned}&&\xi\in[0,L],\ t\geq 0,\\ &\ell-L=\int_{0}^{L}(s^{\prime}-1)\,d\xi,&&t\geq 0,\end{aligned}\right$.
with periodic boundary conditions and appropriate initial data. The 1D finite strain is defined by (A.15), completing the mechanochemical coupling.
A.2. Inifinitesimal strain
To facilitate analytical treatment of the mechanochemical model, we invoke the infinitesimal strain approximation. This is justified biologically, as Hydra cells experience only small-amplitude distortion during mechanochemical signalling, so that the local deformation gradients remain close to the identity [41]. Under the assumption that $|s^{\prime}-1|\ll 1,$ the finite strain $\tilde{\varepsilon}=\tfrac{1}{2}(s^{\prime 2}-1)$ can be replaced by its linearized counterpart
(A.19) $\hat{\varepsilon}(t,\xi)=s^{\prime}(t,\xi)-1\approx\tilde{\varepsilon}(t,\xi),$
[…]
(A.21) $\left\{\begin{aligned} &\begin{aligned} \partial_{t}u&=\kappa\,\varepsilon-\alpha u+D\,\partial_{\xi\xi}u,\\ 0&=\partial_{\xi}\big(E(u)\,\varepsilon\big),\end{aligned}&&\xi\in[0,L],\quad t\geq 0,\\ &l-L=\int_{0}^{L}\varepsilon\,d\xi,&&t\geq 0.\end{aligned}\right$.
Appendix B Well-posedness of the model
@@
(B.1) $u(t,\cdot)=e^{D\Delta t}u_{0}+\int_{0}^{t}e^{D\Delta(t-s)}\left(\kappa\frac{e^{u(s,\cdot)}}{\int_{0}^{1}e^{u(s,y)}dy}-u(s,\cdot)\right)ds\;.$
@@
(B.2) $F(v)=e^{D\Delta t}u_{0}+\int_{0}^{t}e^{D\Delta(t-s)}\left(\kappa\frac{e^{v(s,\cdot)}}{\int_{0}^{1}e^{v(s,y)}dy}-v(s,\cdot)\right)ds.$
@@
(B.3) $V=\left\{v\in X:\|v\|_{X}\leqslant 3\delta\right\}\quad\text{where}\quad\delta=\|u_{0}\|_{W^{1,2}}.$
@@
(B.4) $\|e^{D\Delta t}v\|_{L^{2}}\leqslant C\|v\|_{L^{2}}\quad\text{and}\quad\|\nabla e^{D\Delta t}v\|_{L^{2}}\leqslant Ct^{-\frac{1}{2}}\|v\|_{L^{2}}\quad\text{for }v\in L^{2}(\mathbb{T}).$
@@
(B.5) $\|F(v)\|_{W^{1,2}}$ $\leqslant\|e^{D\Delta t}u_{0}\|_{W^{1,2}}+C\int_{0}^{t}\left\|e^{D\Delta(t-s)}\left(\kappa\frac{e^{v(s,\cdot)}}{\int_{0}^{1}e^{v(s,y)}dy}-v(s,\cdot)\right)\right\|_{W^{1,2}}ds$
@@
(B.6) $\|F(v)\|_{W^{1,2}}$ $\leqslant\|e^{D\Delta t}u_{0}\|_{W^{1,2}}+C\int_{0}^{t}\left(1+(t-s)^{-\frac{1}{2}}\right)\left(\kappa\frac{e^{\|v(s,\cdot)\|_{L^{\infty}}}}{e^{-\|v(s,\cdot)\|_{L^{\infty}}}}+\|v(s,\cdot)\|_{2}\right)ds$
@@
(B.7) $\|F(v)$ $-F(w)\|_{W^{1,2}}\leqslant\int_{0}^{t}\left(1+(t-s)^{-\frac{1}{2}}\right)\left\|\kappa\frac{e^{v}}{\int_{0}^{1}e^{v}dy}-\kappa\frac{e^{w}}{\int_{0}^{1}e^{w}dy}-v+w\right\|_{L^{2}}ds.$
@@
(B.8) $\|F(v)-F(w)\|_{W^{1,2}}$ $\leqslant C(t)\|v-w\|_{X}+C(t)\kappa\sup_{t\in[0,\tau]}\left\|\kappa\frac{e^{v}-e^{w}}{\int_{0}^{1}e^{v}dy}+\kappa e^{w}\frac{\int_{0}^{1}e^{w}dy-\int_{0}^{1}e^{v}dy}{\int_{0}^{1}e^{v}dy\int_{0}^{1}e^{w}dy}\right\|_{L^{2}}.$
@@
(B.9) $e^{w}\left|\int_{0}^{1}(e^{w}-e^{v})dy\right|\big/\left(\int_{0}^{1}e^{v}dy\int_{0}^{1}e^{w}dy\right)\leqslant e^{3C\delta}\int_{0}^{1}|e^{v}-e^{w}|dy$
@@
(B.11) $\|F(v)-F(w)\|_{W^{1,2}}$ $\leqslant C(t)\|v-w\|_{X}+C(t)\kappa e^{C\delta}\sup_{t\in[0,\tau]}\left\|{e^{v}-e^{w}}\right\|_{L^{2}}+C(t)\kappa e^{3C\delta}\sup_{t\in[0,\tau]}\left\|{e^{v}-e^{w}}\right\|_{L^{\infty}}.$
@@
(B.12) $\|F(v)-F(w)\|_{W^{1,2}}\leqslant C(t)\|v-w\|_{X}+C(t)\sup_{t\in[0,\tau]}\left\|{v-w}\right\|_{W^{1,2}}\leqslant C(\tau)\|v-w\|_{X}.$
@@
(B.13) $\frac{d}{dt}\mathcal{J}\big(u(\cdot,t)\big)=D\mathcal{J}\big(u(\cdot,t)\big)u_{t}(\cdot,t)=-\|u_{t}(\cdot,t)\|_{L^{2}}^{2}\leqslant 0\;,$
@@
(B.14) $\frac{D}{2}\int_{0}^{1}\big|u_{x}(x,t)\big|^{2}dx+\frac{1}{2}\int_{0}^{1}\big|u(x,t)\big|^{2}dx$ $\leqslant\mathcal{J}\big(u(\cdot,t)\big)+\kappa\log\left(\int_{0}^{1}e^{u(x,t)}dx\right)$
@@
(B.15) $\int_{0}^{t}\|u_{t}(\cdot,t)\|_{L^{2}}^{2}$ $=J(u_{0})-J\big(u(\cdot,t)\big)\leqslant J(u_{0})+C\|u(\cdot,t)\|_{W{1,2}}+\kappa C\|u(\cdot,t)\|_{L^{\infty}}\leqslant C\;,$
@@
(B.16) $\int_{0}^{t}\|\Delta u(\cdot,t)\|_{L^{2}}^{2}$ $=\int_{0}^{t}\left\|u_{t}(\cdot,t)+u(\cdot,t)+\kappa\frac{e^{u(\cdot,t)}}{\int_{0}^{1}e^{u(y,t)}dy}\right\|_{L^{2}}^{2}$
@@
(B.17) $u\in C\big([0,\infty),W^{1,2}(\mathbb{T})\big),\qquad u\in L^{2}\big([0,\infty),L^{2}(\mathbb{T})\big),\qquad u\in L^{2}\big((0,T),W^{2,2}(\mathbb{T})\big),$
@@
Appendix C Computations of the derivatives
@@
(C.1) $G_{\chi}(\chi,\alpha)(h)$ $=Dh_{xx}-h+(\kappa_{n}+\alpha)\left[\frac{e^{\chi}}{\int e^{\chi}dy}h-\frac{e^{\chi}}{(\int e^{\chi}dy)^{2}}\int e^{\chi}hdy\right],$